Find the interval of convergence of the power series

∞

Ξ£ [(βˆ’1)^𝑛 (𝑛 βˆ’ 2) / 𝑛^2 . 2^𝑛] (π‘₯ βˆ’ 2)^𝑛

𝑛=3


1
Expert's answer
2021-06-16T06:25:31-0400
"\displaystyle\sum_{n=3}^{\infin}\dfrac{(-1)^n(n-2)}{n^22^n}(x-2)^n"

"\bigg|\dfrac{a_{n+1}}{a_n}\bigg|=\bigg|\dfrac{\dfrac{(n+1-2)}{(n+1)^22^{n+1}}(x-2)^{n+1}}{\dfrac{(n-2)}{n^22^n}(x-2)^n}\bigg|"

"=\dfrac{(n-1)n^2}{2(n-2)(n+1)^2}|x-2|\to\dfrac{1}{2}|x-2|\ as\ n\to\infin"


"\dfrac{1}{2}|x-2|<1=>|x-2|<2"

Radius of convergence: "2."


"x=0"


"\displaystyle\sum_{n=3}^{\infin}\dfrac{(-1)^n(n-2)}{n^22^n}(0-2)^n=\displaystyle\sum_{n=3}^{\infin}\dfrac{n-2}{n^2}"

Use the Limit Comparison Test with


"\displaystyle\sum_{n=3}^{\infin}b_n=\displaystyle\sum_{n=3}^{\infin}\dfrac{1}{n}"

"\lim\limits_{n\to\infin}\dfrac{\dfrac{n-2}{n^2}}{\dfrac{1}{n}}=1>0"

The garmonic series "\displaystyle\sum_{n=3}^{\infin}\dfrac{1}{n}" diverges as "p" series with "p=1."

Then the series "\displaystyle\sum_{n=3}^{\infin}\dfrac{n-2}{n^2}" diverges by the Limit Comparison Test.


"x=4"


"\displaystyle\sum_{n=3}^{\infin}\dfrac{(-1)^n(n-2)}{n^22^n}(4-2)^n=\displaystyle\sum_{n=3}^{\infin}\dfrac{(-1)^n(n-2)}{n^2}"

Use the Alternating Series Test


"a_n=\dfrac{n-2}{n^2}"

"\lim\limits_{n\to\infin}a_n=\lim\limits_{n\to\infin}\dfrac{n-2}{n^2}=0"

"a_{n+1}=\dfrac{n+1-2}{(n+1)^2}<\dfrac{n-2}{n^2}=a_n, n\geq 3"

Then the series "\displaystyle\sum_{n=3}^{\infin}\dfrac{(-1)^n(n-2)}{n^2}" converges by the Alternating Series Test.

Therefore the interval of convergence is "(0, 4]."